# Debiasing astro-Photometric Observations with Corrections Using Statistics (DePhOCUS)

### v1

@@ Acknowledgements
This work makes use of the following publicly available codes: Pandas (Pandas development team, 2020; Wes McKinney, 2010), Matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), Astropy (Astropy Collaboration et al., 2013, 2018), Astroquery (Ginsburg et al., 2019).
@@ Appendix A Correction files
Correction system based on NEOCC, aB, aBC and aBCO corrections with confidence level above a certain $p$-value can be found as separate data files, named “corr_DePhOCUS_p[XXX].txt”, with $p=0.90$ being the optimum correction according to the results.
@@ Appendix B Additional Figures
(Part 2 - continued on next page)
Figure 16: Results from aBCO bias analysis with data from 394 out of 468 asteroids with the average bias and its standard error, sorted in decreasing order of the number of observations (only showing results with more than 10 observations). The horizontal axis lists the combination of the catalog (one letter format) and the three-letter MPC observatory code. The results are separated by color band with according band bias (colored line) including error margins (blue area) from the aB analysis in Tab. 2 in comparison, leaving out the following bands with just one observatory: Y, H, K (only W91) and y (only F51)

### v2

@@ Acknowledgements
This work makes use of the following publicly available codes: Pandas (Pandas development team, 2020; Wes McKinney, 2010), Matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), Astropy (Astropy Collaboration et al., 2013, 2018), Astroquery (Ginsburg et al., 2019), diptest (Hartigan, 1985).
@@ Appendix A Correction files
Correction system based on NEOCC, aB, aBC and aBCO corrections with confidence level $\gamma$ can be found as separate data files, named “corr_DePhOCUS_g[XXX].txt”, with $\gamma=0.90$ being the optimum correction according to the results.
@@ Appendix B Additional Figures
Figure 13: Histogram for the number of observations for each of the 468 asteroids in full range (left) and cut-out up to 1000 observations (right).
@@
Figure 16: Results from aBCO bias analysis with data from 394 out of 468 asteroids with the average bias and its standard error, sorted in decreasing order of the number of observations (only showing results with more than 50 observations). The horizontal axis lists the combination of the catalog (one letter format, ’-’ meaning not specified) and the three-letter MPC observatory code. The results are separated by color band with according band bias (colored line) including error margins (blue area) from the aB analysis in Tab. 2 in comparison, leaving out the following bands with just one observatory: H, J, K, Y (only W91), y (only F51), I (only L10), u (only 309) and B (only 270)